\begin{thebibliography}{8}
%\bibitem{Two pages} C. Apostol, L.A. Fialkow, D.A. Herrero, D. Voiculescu, \textit{Approximation of Hilbert Space Operators II}, Research Notes in Mathematics, 102, Pitman (Advanced Publishing Program), Boston, 1984.
\bibitem{joint} H. Baklouti, K. Feki, On joint spectral radius of commuting operators in Hilbert spaces, \textit{Linear Algebra Appl.} 557 (2018) 455-463.
\bibitem{Bhatia}
R. Bhatia, P. $\check{\text S}$emrl, Orthogonality of matrices and some distance problems, \textit{Linear Algebra Appl.} 287 (1999), 77--85.
\bibitem{2013} T. Bhattacharyya, P. Grover, Characterization of Birkhoff-James orthogonality, \textit{J. Math. Anal. Appl.} 407 (2013), 350--358.
\bibitem{Davidson} K. R. Davidson, {\it Nest algebras, Triangular forms for operator algebras on Hilbert space,} {Pitman Research Notes in Mathematics Series, 191}, New York, 1988.
\bibitem{2017} P. Grover, Orthogonality of matrices in the Ky Fan $k$-norms, \textit{Linear Multilinear Algebra} 65 (2017), 496--509.
\bibitem{2014} P. Grover, Orthogonality to matrix subspaces, and a distance formula, \textit{Linear Algebra Appl.} 445 (2014), 280--288.
%\bibitem{thesis} P. Grover, \textit{Some Problems in Differential and Subdifferential Calculus of Matrices}, Ph.D. thesis, Indian Statistical Institute, India, 2014.

\bibitem{2019} P. Grover, S. Singla,  Best approximations, distance formulas and orthogonality in $C^*$-algebras, {\it J. Ramanujan Math. Soc.} 36 (2021), 85–91.
\bibitem{2020} P. Grover, S. Singla, \textit{Birkhoff-James orthogonality and applications : A survey.} Operator Theory, Functional Analysis and Applications, eds. M. A. Bastos, L. Castro, A. Y. Karlovich, Oper. Theory Adv. Appl., Birkh\"{a}user Cham, 282 (2021), 293--315.
%\bibitem{2021} P. Grover, S. Singla,  A distance formula  for tuples of operators, {\it communicated}. 

\bibitem{hiriart} J.B. Hiriart-Urruty, C. Lemar$\grave{\text e}$chal, \textit{Fundamentals of Convex Analysis}, Springer, 2000.

\bibitem{magajna} B. Magajna, On the distance to finite-dimensional subspaces in operator algebras, {\it J. London Math. Soc. (2)} 47 (1993), 516--532.

\bibitem{arxiv} A. Mal, K. Paul, Birkhoff-James orthogonality to a subspace of operators defined between Banach spaces, {\it J. Operator Theory} 85 (2021), 463--474.

\bibitem{radiusparallel} M. Mehrazin, M. Amyari, A. Zamani,  Numerical radius parallelism of Hilbert space operators, {\it Bull. Iranian Math. Soc.} 46 (2020), 821--829.

\bibitem{normparallel} M. S. Moslehian, A. Zamani, Norm-parallelism in the geometry of Hilbert $C^*$-modules, {\it Indag. Math.} 27 (2016) 266--281.
 
%\bibitem{li} C.K. Li, H. Schneider, Orthogonality of Matrices, \textit{Linear Algebra Appl.} 347 (2002) 115-122.

%\bibitem{Radii} K. Paul, Translatable radii of an operator in the direction of another operator, \textit{Sci. Math.}, 2 (1999), 119--122.

\bibitem{bilinear} S. Roy, T. Senapati, D. Sain, Orthogonality of bilinear forms and application to matrices, \textit{Linear Algebra Appl.} 615 (2021), 104--111.

%\bibitem{marshall} A.W. Marshall, I. Olkin, \textit{Inequalities: Theory of Majorization and Its Applications}, Academic Press, 1979.
%\bibitem{rockafellar} R.T. Rockafellar, \textit{Convex Analysis}, Princeton University Press, 1972.
\bibitem{Singer} I. Singer, \textit{Best Approximation in Normed Linear Spaces by Elements of Linear Subspaces}, Springer, 1970.
\bibitem{new} S. Singla, Gateaux derivative of $C^*$ norm, \textit{Linear Algebra Appl.} 629 (2021), 208-218.
%\bibitem{derivation} J. Stampfli, The norm of a derivation, \textit{Pac. J. Appl. Math.} 33 (1970), 737--747.
%\bibitem{watson1} G. A. Watson, \textit{Approximation Theory and Numerical Methods}, John Wiley \& Sons, 1980.
\bibitem{Watson} G. A. Watson, Characterization of the subdifferential of some matrix norms, \textit{Linear Algebra Appl.} 170 (1992), 33--45.
\bibitem{watson93} G. A. Watson, On matrix approximation problems with Ky Fan $k$ norms, \textit{Numer. Algorithms} 5 (1993),  263--272.
\bibitem{Wojcik} P. W$\acute{\text o}$jcik, Birkhoff orthogonality in classical M-ideals, {\it J. Aust. Math. Soc.} 103 (2017), 279--288.
%\bibitem{LO} J. Weidmann, \textit{Linear Operators in Hilbert Spaces}, Springer, New York, 1980.
\bibitem{Zietak2}  K. Zi\k{e}tak, From the strict Chebyshev approximant of a vector to the strict spectral approximant of a matrix, {\it Banach Center Publ., 112} {\it Polish Acad. Sci. Inst. Math.} Warsaw, 2017.
\bibitem{Zietak} K. Zi\k{e}tak, Subdifferentials, faces, and dual matrices, \textit{Linear Algebra Appl.}, 185 (1993), 125--141.
\bibitem{convex} C. Z\u{a}linescu, {\it Convex analysis in general vector spaces}, World Scientific Publishing Co., Inc., River Edge, NJ, 2002.
\end{thebibliography}

\documentclass[a4paper,12pt]{amsart}
\usepackage{amsmath,latexsym,amsfonts,amssymb,mathrsfs,amsthm,hyperref, enumitem}
\usepackage[T1]{fontenc}

\title{Subdifferential of the joint numerical radius}
\author[Grover]{Priyanka Grover}
\author[Singla]{Sushil Singla}
\address{Department of Mathematics, Shiv Nadar University, NH-91, Tehsil Dadri, Gautam Buddha Nagar, U.P. 201314, India.}
\email{priyanka.grover@snu.edu.in, ss774@snu.edu.in}
\subjclass[2010]{15A60, 58C20, 47A12}
\keywords{Subdifferential set, Gateuax derivative, Joint numerical range, trace class, Birkhoff-James orthogonality}

\newcommand{\C}{\mathbb{C}}
\newcommand{\R}{\mathbb{R}}
\newcommand{\tr}{\mathop{{\rm trace}}}
\newcommand{\conv}{\mathop{{\rm co}}}
\newcommand{\h}{\mathop{{\rm Re}}}

\newcommand{\hc}{{\mathcal H}}
\newcommand{\hk}{{\mathcal K}}

\newcommand{\dist}{{\mathop{\rm dist}}}

\newtheorem{theorem}{Theorem}

\theoremstyle{definition}

\newtheorem{corollary}{Corollary}

\newtheorem{prop}{Proposition}

\newcommand{\Hil}{\mathcal{H}}

\newcommand{\tens}[1]{\mathbin{\mathop{\otimes}\limits_{#1}}}
\newcommand{\bra}{\langle}
\newcommand{\ket}{\rangle}

\begin{document}
\begin{abstract}
An expression for the subdifferential of the joint numerical radius is obtained. Its applications to the best approximation problems in the joint numerical radius are discussed.

\end{abstract}

\setlength{\parindent}{0pt}
\setlength{\parskip}{1.6ex}

\maketitle

\numberwithin{theorem}{section}
\numberwithin{corollary}{section}
\numberwithin{equation}{section}
\numberwithin{defn}{section}
\numberwithin{prop}{section}

\section{Introduction}
Let $\mathbb M_{p,q}(\mathbb C)$ be the set of $p\times q$ matrices over $\mathbb C$ with a given norm. Let $f: \mathbb M_{p,q}(\mathbb C)\rightarrow \R$ be a continuous convex function. Let $A \in \mathbb M_{p,q}(\mathbb C)$. The \emph{subdifferential} of $f$ at $A$, denoted by $\partial f(A)$, is defined as
$$\partial f(A)=\{C\in\mathbb M_{p,q}(\mathbb C): f(B)\geq f(A)+\h\tr((B-A)^*C) \text{ for all } B\in\mathbb M_{p,q}(\mathbb C)\}.$$
The right hand derivative of $f$ and the subdifferential of $f$ are related as follows. For $B\in\mathbb M_{p,q}(\mathbb C)$,
 \begin{equation}\label{chap5:7898}\lim\limits_{t\rightarrow 0^+} \dfrac{f(A+tB) - f(A)}{t} = \max\{\h\tr(C^*A) : C\in \partial f(A)\}.\end{equation} 
Characterizations of subdifferentials of matrix norms has been of interest to many mathematicians. Let $\mathbb M_n(\mathbb C)$ be the set of $n\times n$ matrices over $\mathbb C$. For $A\in\mathbb M_n(\mathbb C)$, let  $s_1(A)\geq \cdots\geq s_n(A)$ be the singular values of $A$. Let $|||\cdot|||$ denote a  {unitarily invariant norm} on $\mathbb M_n(\mathbb C)$ (that is, for any unitary matrices $U$ and $U'$, we have $|||UAU'|||=|||A|||$). Then there is a unique {symmetric gauge function} $\Phi$ on $\mathbb R^n$ such that $|||A|||=\Phi((s_1(A), \ldots, s_n(A))$ for every $A\in \mathbb M_n(\mathbb C)$. In \cite[Theorem 3.1, Theorem 3.2]{Zietak}, it was shown that for $A\in\mathbb M_n(\mathbb C)$,
\begin{align*}\partial |||A||| = \{& U \text{diag}(d_1,\ldots,d_n) U'^*: A=U \Sigma U'^* \text{ is a singular value decomposition  of A,}\\
& \sum s_i(A) d_i=|||A|||=\Phi((s_1,\ldots, s_n)), \Phi^*((d_1,\ldots, d_n))=1\}.\end{align*}
This was an improvement of Theorem 2 of \cite{Watson}, where an expression of $\partial |||\cdot|||$ was given in  $\mathbb M_n(\mathbb R)$. 
In \cite[Theorem 1]{watson93}, the above result was proved using a different approach. Let $\|\cdot\|$ be the \emph{operator norm} (or the \emph{spectral norm}) on $\mathbb M_n(\mathbb C)$, defined as : $$\|A\|=\max\{\|Au\|: \|u\|=1\}.$$ The operator norm is a unitarily invariant norm and we have the following. For $A\in M_n(\mathbb C)$, 
$$\partial \|A\|=\conv\{uv^*:\|u\|=\|v\|=1, Av=\|A\|u\},$$ where $\conv(S)$ denotes the convex hull of a set $S$. For $1\leq k\leq n$, the Ky Fan $k$-norm $\|\cdot\|_{(k)}$ is defined as $$\|A\|_{(k)}=s_1(A)+\dots +s_k(A).$$
The subdifferential set of the {Ky Fan $k$-norms} on $\mathbb M_n(\mathbb C)$ was obtained in Theorem 2.7 of \cite{2017}. Another useful norm on $\mathbb M_n(\mathbb C)$ is the \emph{numerical radius}, defined as
$$w(A)=\max_{\|x\|=1} \left|\langle x, Ax\rangle\right|.$$ More generally, we consider the \emph{joint numerical radius} of a tuple of matrices defined as follows.
Let $A_1, \dots, A_d\in \mathbb M_n(\mathbb C)$. Let $\boldsymbol{A}=(A_1, \dots, A_d):\mathbb C^n \rightarrow (\mathbb C^n)^d$ be defined as $\boldsymbol{A}x=(A_1x, \dots, A_dx)$ for all $x\in\mathbb C^n$. The {joint numerical radius} of $\boldsymbol{A}$ is defined as 
$$\omega(\boldsymbol{A})=\max_{x\in \mathbb C^n,\ \|x\|=1}\left(\sum\limits_{k=1}^d\big|\langle x| A_kx\rangle\big|^2\right)^{1/2}.$$
%Note that the joint numerical radius $\omega(\cdot)$ is a norm on $\mathbb M_n(\mathbb C)^d$.  
%The main results of the paper are given below. Few notations are in order. 
%For a subset $S$ of $\mathbb M_{p,q}(\mathbb C)$, let $\conv(S)$ denote the closed convex hull of $S$. 
For $x\in\mathbb C^n$, let $x\bar{\tens{}} x$ be the rank one operator on $\mathbb C^n$ defined as $x\bar{\tens{}} x(y) = \bra y | x\ket x$ for all $y\in\C^n$. We will use the same symbol $x\bar{\tens{}} x$ for the rank one operator as well as its matrix representation.  Let $\boldsymbol{0}=(0,\dots, 0)\in\mathbb M_n(\C)^d$. The main result of this paper is as follows.
\begin{theorem}\label{main}
Let $\boldsymbol{A}\in\mathbb M_n(\C)^d\setminus\{\boldsymbol{0}\}$. Then
\begin{enumerate}[label=(\alph*)]
\item the subdifferential of $\omega(\cdot)$ at $\boldsymbol{A}$ is given by
\begin{align}
\partial \omega(\boldsymbol{A})=\conv\bigg\{&\frac{1}{\omega(\boldsymbol{A})} \left(\overline{\langle x|A_1x\rangle} x\bar{\tens{}}x, \dots, \overline{\langle x|A_dx\rangle} x\bar{\tens{}}x\right)  :  \|x\|=1,\nonumber\\
&\omega(\boldsymbol{A})=\left(\sum\limits_{k=1}^d\big|\langle x|A_kx\rangle\big|^2\right)^{1/2}\bigg\},\label{subdifferential}
\end{align}
and
\item for $\boldsymbol{B}\in\mathbb M_n(\C)^d$,
\begin{equation*}\lim\limits_{t\rightarrow 0^+}\frac{\omega(\boldsymbol{A}+t\boldsymbol{B}) -\omega(\boldsymbol{A})}{t} = \frac{1}{\omega(\boldsymbol{A})}\max_{ \|x\|=1,\ \omega(\boldsymbol{A})=\left(\sum\limits_{k=1}^d|\langle x|A_kx\rangle|^2\right)^{1/2}}\h\sum\limits_{k=1}^d\langle x|A_kx\rangle\overline{\langle x|B_kx\rangle}.\end{equation*}
\end{enumerate}
\end{theorem}

For $\boldsymbol{\lambda}=(\lambda_1, \dots, \lambda_d)\in\mathbb C^d$ and $\boldsymbol{B}=(B_1,  \dots, B_d)\in\mathbb M_n(\mathbb C)^d$, let $\boldsymbol{\lambda B} = (\lambda_1 B_1, \dots, \lambda_d B_d)$. As a consequence of Theorem \ref{main}, we obtain the following result.
\begin{corollary}\label{101010}
Let $\boldsymbol{A}=(A_1, \dots, A_d)$, $\boldsymbol{B}=(B_1,  \dots, B_d)\in\mathbb M_n(\mathbb C)^d$. Then 
\begin{equation}\label{1111}
\omega(\boldsymbol{A}+\boldsymbol{\lambda B}) \geq \omega(\boldsymbol{A}) \text{ for all } \boldsymbol{\lambda}\in \C^d
\end{equation}
if and only if there exist $h$ unit vectors $x_1,\ldots,x_h \in \C^n$ with $\omega(\boldsymbol{A})=\left(\sum\limits_{k=1}^d\big|\langle x_i|A_kx_i\rangle\big|^2\right)^{1/2}$ for all $1\leq i\leq h$ and there exist $h$ positive numbers $t_1,\ldots,t_h >0$ with $t_1+\cdots+t_h=1$ such that $\sum\limits_{i=1}^h t_i\langle x_i| A_kx_i\rangle\overline{\langle x_i|B_kx_i\rangle} =0$ for all $1\leq k\leq d$.
\end{corollary}

When $d=1$, the sufficiency of the above condition was given in \cite[Theorem 2.11]{arxiv}. In Section \ref{proofs}, we give proofs of Theorem \ref{main} and Corollary \ref{101010}. We also obtain analogous results for the  \emph{joint operator norm}. Finally, we end with some remarks in Section  \ref{Remarks}.

\section{Proofs}\label{proofs}
To prove Theorem \ref{main}, we will need the following propositions from the subdifferential calculus. 

%Let $X$ and $Y$ be Banach spaces. Let $\mathscr B(X, Y)$  be the space of bounded $\C$-linear operators from $X$ to $Y$ with operator norm.

\begin{prop}\label{chainrule} Let $T_1: \mathbb M_{p,q}(\mathbb C)\rightarrow \mathbb M_{r,s}(\mathbb C)$ be a linear map. Let $B\in \mathbb M_{r,s}(\mathbb C)$. Let $T_2: \mathbb M_{p,q}(\mathbb C)\rightarrow \mathbb M_{r,s}(\mathbb C)$ be the  affine map defined as $T_2(A)=T_1(A)+ B$. Let $g: \mathbb M_{r,s}(\mathbb C) \rightarrow \R$ be a continuous convex function. Then for $A \in \mathbb M_{p,q}(\mathbb C)$,
\begin{equation*}
\partial (g\circ T_2)(A)=T_1^* \partial g(T_2(A)).\label{1.2}
\end{equation*}
\end{prop}

\begin{prop}\label{maxsubdifferential}
Let $J$ be a compact set in some metric space. Let $\{f_j\}_{j\in J}$ be a collection of continuous convex functions from $\mathbb M_{p,q}(\mathbb C)$ to $\R$ such that for $A\in \mathbb M_{p,q}(\mathbb C)$, the maps $j\rightarrow f_j(A)$ are upper semi-continuous. Let $f: \mathbb M_{p,q}(\mathbb C)\rightarrow \R$ be defined as $f(A)=\sup\{f_j(A):j\in J\}$. Let $J(A)=\{j \in J: f_j(A)=f(A)\}$. Then
\begin{equation*}
\partial f(A)=\conv\left(\cup\left\{\partial f_j(A): j\in J(A)\right\}\right).\label{maxsubdiff}
\end{equation*}
\end{prop}

The proofs of these can be found in Theorem 4.2.1 and Theorem 4.4.2 of  \cite{hiriart}. In this book the author deals with real valued convex functions on Euclidean space $\mathbb R^n$. The same proofs can be extended to real valued continuous convex functions on a normed space also (see \cite{convex} for more detail). Now we prove Theorem \ref{main}.

\textit{Proof of Theorem \ref{main}.} 
\begin{enumerate}[label=(\alph*)]
\item In \cite{joint}, it was shown that $\omega(\boldsymbol{A})$ can also be expressed as \begin{equation}\omega(\boldsymbol{A})=\max_{x\in \C^n,\ \|x\|=1}\max_{(\lambda_1, \dots, \lambda_d)\in\C^d, \|(\lambda_1,  \dots, \lambda_d)\|=1}\left| \sum\limits_{k=1}^d \lambda_k\langle x|A_kx\rangle\right|.\label{synk1}\end{equation} 
Let $\boldsymbol{\lambda} = (\lambda_1, \dots, \lambda_d)\in\C^d$ and let $x\in\C^n$. Let $\boldsymbol{C}=(C_1, \dots, C_d)\in\mathbb M_n(\C)^d.$ Let $T_{x,\boldsymbol{\lambda}}: \mathbb M_n(\C)^d\rightarrow \C$ be the linear map defined as
$$T_{x,\boldsymbol{\lambda}}(\boldsymbol{C})=\sum\limits_{k=1}^d \lambda_k\langle x| C_kx\rangle.$$ Let $z\in\C$. Let $g:\C\rightarrow\R$ be the map defined as $g(z)=|z|$.
Let $f_{x,\boldsymbol{\lambda}}: \mathbb M_n(\C)^d \rightarrow \R$ be the map defined as  $f_{x,\boldsymbol{\lambda}}=g$ $\circ\ T_{x,\boldsymbol{\lambda}}$. %We note that for $\boldsymbol A\neq \boldsymbol{0}$, $f_{x,\boldsymbol{\lambda}}$ is a differentiable map. Hence, $\partial f_{x,\boldsymbol{\lambda}}(\boldsymbol A)$ is the singleton set containing the derivative $\text{D} f_{x,\boldsymbol{\lambda}}(\boldsymbol A)$.  
Let $J$ be the compact set $\left\{(x,\boldsymbol{\lambda})\in \C^n\times\C^d :\ \|x\|=1, \|\boldsymbol{\lambda}\|=1\right\}$. Note that for $\boldsymbol{C}\in \mathbb M_n(\C)^d$, the map $(x,\boldsymbol{\lambda})\rightarrow f_{x,\boldsymbol{\lambda}}(\boldsymbol{C})$ is continuous.  Now \eqref{synk1} can be rewritten as
$$\omega(\boldsymbol{A})=\max\{f_{x,\boldsymbol{\lambda}}(\boldsymbol{A}): (x,\boldsymbol{\lambda})\in J\}.$$ 
Let $J(\boldsymbol{A})=\{(x,\boldsymbol{\lambda})\in J: f_{x,\boldsymbol{\lambda}}(\boldsymbol{A})=\omega(\boldsymbol{A})\}.$ 
By Proposition \ref{maxsubdifferential}, 
$$\partial \omega(\boldsymbol{A})=\conv\left(\cup\left\{\partial f_{x,\boldsymbol{\lambda}}(\boldsymbol{A}): (x,\boldsymbol{\lambda})\in J(\boldsymbol{A})\right\}\right).$$
Let $(x,\boldsymbol{\lambda})\in J(\boldsymbol{A})$. Then  $\sum\limits_{k=1}^d \lambda_k\langle x|A_kx\rangle\neq 0$. By Proposition \ref{chainrule}, we get
\begin{eqnarray*}
\partial f_{x,\boldsymbol{\lambda}}(\boldsymbol{A}) &=& T_{x,\boldsymbol{\lambda}}^*\partial g(T_{x,\boldsymbol{\lambda}}(\boldsymbol{A}))\\
&=& T_{x,\boldsymbol{\lambda}}^*\partial g\bigg(\sum\limits_{k=1}^d \lambda_k\langle x| A_kx\rangle\bigg)\\
&=& \left\{T_{x,\boldsymbol{\lambda}}^*\left( \frac{\sum\limits_{k=1}^d \lambda_k\langle x| A_kx\rangle}{\left|\sum\limits_{k=1}^d \lambda_k\langle x|A_kx\rangle\right|}\right)\right\}.
\end{eqnarray*}

Now $T_{x,\boldsymbol{\lambda}}^*: \C\rightarrow \mathbb M_n(\C)^d$ is the unique map satisfying 
\begin{equation}\label{chap5:11111}\tr [(T_{x,\boldsymbol{\lambda}}^*(z))^* \boldsymbol{C}]=\overline{z} \ T_{x,\boldsymbol{\lambda}}(\boldsymbol{C}).\end{equation}

If $T_{x,\boldsymbol{\lambda}}^*(z) = (T_1, \dots, T_d)$, then \eqref{chap5:11111} gives $$\sum\limits_{k=1}^d\tr(T_k^*C_k) = \sum\limits_{k=1}^d\overline{z}\lambda_k\langle x|C_k x\rangle.$$

This implies that for $z\in\C$, $T_{x,\boldsymbol{\lambda}}^*(z)=z\left(\overline{\lambda_1} x\bar{\tens{}}x, \overline{\lambda_2} x\bar{\tens{}}x, \dots, \overline{\lambda_d} x\bar{\tens{}}x\right)$.
So
$$\partial f_{x,\boldsymbol{\lambda}}(\boldsymbol{A})=\left\{ \frac{\sum\limits_{k=1}^d \lambda_k\langle x| A_kx\rangle}{\left|\sum\limits_{k=1}^d \lambda_k\langle x| A_kx\rangle\right|}\left(\overline{\lambda_1} x \bar{\tens{}}x, \overline{\lambda_2} x\bar{\tens{}} x, \dots, \overline{\lambda_d} x \bar{\tens{}}x\right)\right\}.$$

This gives
\begin{equation}\partial \omega(\boldsymbol{A})=\conv\left\{\frac{\sum\limits_{k=1}^d \lambda_k\langle x|A_kx\rangle}{\omega(\boldsymbol{A})}(\overline{\lambda_1} x\bar{\tens{}} x, \dots, \overline{\lambda_d} x\bar{\tens{}} x): (x,\boldsymbol{\lambda})\in J(\boldsymbol{A})\right\}.\label{chap5:75757}\end{equation}

For each $(x,\boldsymbol{\lambda})\in J(\boldsymbol{A})$, we have $$\left(\sum\limits_{k=1}^d|\langle x|A_kx\rangle|^2\right)^{1/2}\leq \omega(\boldsymbol{A}) = \left|\sum\limits_{k=1}^d \lambda_k\langle x|A_kx\rangle\right|\leq\left(\sum\limits_{k=1}^d|\langle x| A_kx\rangle|^2\right)^{1/2}.$$

The last inequality follows by the Cauchy-Schwarz inequality. Hence $$\left|\sum\limits_{k=1}^d \lambda_k\langle x| A_kx\rangle\right| = \left(\sum\limits_{k=1}^d\big|\langle x| A_kx\rangle\big|^2\right)^{1/2}\left(\sum\limits_{k=1}^d|\lambda_k|^2\right)^{1/2}.$$ By the condition of equality in the Cauchy-Schwarz inequality, there exists $\alpha\in\C$ such that $\left(\overline{\lambda_1}, \dots, \overline{\lambda_d}\right) = \alpha \left(\langle x| A_1x\rangle, \dots, \langle x| A_dx\rangle\right).$ This gives $\alpha = \left(\sum\limits_{k=1}^d|\langle x | A_kx\rangle|^2\right)^{-1/2}$. Substituting the value of $\overline{\lambda_k}$ in \eqref{chap5:75757}, we get \eqref{subdifferential}.
\item This follows as an application of \eqref{chap5:7898} and \eqref{subdifferential}.
\end{enumerate}\qed

%The right hand derivative of a continuous convex function $f: X\rightarrow \R$ and the subdifferential of $f$ are related as follows.
 %\begin{equation*}\lim\limits_{t\rightarrow 0^+} \dfrac{f(x+ty) - f(x)}{t} = \max\{\text{Re }v^*(w) : v^*\in \partial f(x)\}.\end{equation*}
%As an application of Theorem \ref{main}, we get the following expression for the right hand derivative of the joint numerical radius. For $\boldsymbol{A}, \boldsymbol{B}\in\mathbb M_n(\C)^d$, we have
%\begin{equation}\label{chap5:7898}\lim\limits_{t\rightarrow 0^+}\frac{\omega(\boldsymbol{A}+t\boldsymbol{B}) -\omega(\boldsymbol{A})}{t} = \frac{1}{\omega(\boldsymbol{A})}\max_{ \|x\|=1,\ \omega(\boldsymbol{A})=(\sum\limits_{k=1}^d|\langle x|A_kx\rangle|^2)^{1/2}}\h\sum_{k=1}^d\langle x|A_kx\rangle\overline{\langle x|B_kx\rangle}.\end{equation}

Using Theorem \ref{main}, we give the proof of Corollary \ref{101010}. The idea is similar to  \cite[Theorem 2.6]{2013} and \cite[Theorem 1]{2014}. 

\textit{Proof of Corollary \ref{101010}.} Without loss of generality, let $\boldsymbol{A}\neq\boldsymbol{0}$. Let $T_1:\C^d\rightarrow \mathbb M_n(\C)^d$ be the linear map defined as $T_1(\boldsymbol{\lambda})=\boldsymbol{\lambda}\boldsymbol{B}$. Let $T_2:\C^d\rightarrow \mathbb M_n(\C)^d$ be defined as the affine map $L(\boldsymbol{\lambda})=T_1(\boldsymbol{\lambda})+\boldsymbol{A}$ for all $\boldsymbol{\lambda}\in\C^d$.
It is easy to see that 
\begin{equation*}
\omega(\boldsymbol{A}+\boldsymbol{\lambda B}) \geq \omega(\boldsymbol{A}) \text{ for all } \boldsymbol{\lambda}\in \C^d\text{ if and only if }\boldsymbol{0}\in \partial(\omega\circ T_2)(0,\dots, 0).
\end{equation*}

By Proposition \ref{chainrule}, we get
\begin{equation}
\omega(\boldsymbol{A}+\boldsymbol{\lambda}\boldsymbol{B}) \geq \omega(\boldsymbol{A}) \text{ for all } \boldsymbol{\lambda}\in \C^d\text{ if and only if }\boldsymbol{0}\in T_1^* \partial \omega(\boldsymbol{A}). \label{extremum}
\end{equation}

The map $T_1^*: \mathbb M_n(\C)^d\rightarrow \C^d$ is given by $T_1^*(\boldsymbol{C}) = \left(\overline{\tr(C_1^*B_1)}, \dots, \overline{\tr(C_d^*B_d)}\right)$ for all $\boldsymbol{C}=(C_1, \dots, C_d)\in \mathbb M_n(\C)^d$. Therefore
\begin{align}T_1^*\partial\omega(\boldsymbol{A})  = \conv\bigg\{& \dfrac{1}{\omega(\boldsymbol{A})}\left(\overline{\langle x|B_1x\rangle} \langle x|A_1x\rangle, \dots, \overline{\langle x|B_dx\rangle} \langle x|A_dx\rangle\right) : \ \|x\|=1,\nonumber\\
&\omega(\boldsymbol{A})=\left(\sum\limits_{k=1}^d\big|\langle x|A_kx\rangle\big|^2\right)^{1/2}\bigg\}.\label{chap5:8989}\end{align}
The result follows by substituting \eqref{chap5:8989} in \eqref{extremum}. \qed

Let $(X,\|\cdot\|)$ be a normed space. An element $x\in X$ is said to be \emph{(Birkhoff-James) orthogonal} to a subspace $W$ in $\|\cdot\|$ if \begin{equation}\|x+y\|\geq \|x\|\text{ for all }y\in W.\label{synn:new}\end{equation} If $W$ is a one-dimensional subspace generated by $z$ and \eqref{synn:new} is satisfied, then we say that $x$ is orthogonal to $z$.  For $(\mathbb M_n(\mathbb C)^d, \omega(\cdot))$, \eqref{1111} is equivalent to saying that $\boldsymbol{A}$ is orthogonal to the subspace $\{\boldsymbol{\lambda B}: \boldsymbol{\lambda}\in\C^d\}$ in $\omega(\cdot)$. In the proof of Corollary \ref{101010}, if we take $T_1:\C\rightarrow \hc^d$ to be the linear map defined as $T_1(\lambda)=\lambda\boldsymbol{B}$  and $T_2:\C\rightarrow \hc^d$ to be the affine map $T_2(\lambda)=T_1(\lambda)+\boldsymbol{A}$, then we get the following characterization of orthogonality in $(\mathbb M_n(\mathbb C)^d, \omega(\cdot))$.

\begin{theorem}\label{100000}
Let $\boldsymbol{A}=(A_1, \dots, A_d), \boldsymbol{B}=(B_1,  \dots, B_d)\in\mathbb M_n(\mathbb C)^d$. Then $\boldsymbol{A}$ is orthogonal to $\boldsymbol{B}$ if and only if there exist $h$ unit vectors $x_1,\ldots,x_h \in \Hil$ with $\omega(\boldsymbol{A})=\left(\sum\limits_{k=1}^d\big|\langle x_i|A_kx_i\rangle\big|^2\right)^{1/2}$ for all $1\leq i\leq h$ and there exist $h$ positive numbers $t_1,\ldots,t_h>0$ with $t_1+\cdots+t_h=1$ such that $
\sum\limits_{k=1}^d \sum\limits_{i=1}^h t_i\langle x_i|A_kx_i\rangle\overline{\langle x_i| B_kx_i\rangle}=0.$
\end{theorem}

The joint operator norm of $\boldsymbol{A}$ is equal to $\sup
\left\{\left(\sum\limits_{k=1}^d\|A_kx\|^2\right)^{1/2} : x\in\mathbb C^n, \|x\|=1\right\}$. For the joint operator norm, an analogous result to Theorem \ref{100000} was proved in \cite[Corollary 3.4]{2013}. 
A bounded linear map $T$ from a finite dimensional space $X$ to a Banach space $Y$ can be identified with the continuous function from the unit sphere $S_X$ of $X$ to $Y$, defined by $\hat{T}(x) = T(x)$ for all $x\in S_X$. Let $C(S_X, Y)$ denote the space of continuous functions from $S_X$ to $Y$ with the supremum norm $\|\cdot\|_{\infty}$. Then we have $\|T\|=\|\hat{T}\|_{\infty}$. In particular, the space $\mathbb M_n(\C)^d$ equipped with the joint operator norm is isometrically isomorphic to a closed subspace of $C(S_{\C^n}, (\C^n)^d)$. In \cite{bilinear}, this identification was used to give an alternate proof of orthogonality to one dimensional subspaces in $M_n(\C)$ given in \cite[Theorem 1]{Bhatia}. 
We use this identification to prove the following result for the joint operator norm, analogous to Corollary \ref{101010}.

\begin{theorem}\label{newproof} Let $\boldsymbol{A}, \boldsymbol{B}\in  \mathbb M_n(\C)^d$. Then \begin{equation}\|\boldsymbol{A+\lambda B}\|\geq \|\boldsymbol{A}\| \text{ for all } \boldsymbol{\lambda}\in \C^d\label{chap3:eqn1}\end{equation} if and only if 
there exist $h$ unit vectors $x_1,\ldots,x_h \in \C^n$ with $\|\boldsymbol{A}x_i\|=\|\boldsymbol{A}\|$ for all $1\leq i\leq h$  and there exist $h$ positive numbers $t_1,\ldots,t_h>0$ with $t_1+\cdots+t_h=1$ such that $\sum\limits_{i=1}^h t_i\langle A_kx_i|B_kx_i \rangle=0 \text{ for all } 1\leq k\leq d.$ 
Moreover, we have $1\leq h\leq 2d+1$.
\end{theorem}
\begin{proof} If $\boldsymbol{A}\in\mathbb C^d\boldsymbol{B}$, then the theorem holds trivially and $\boldsymbol{A}=0$ if it satisfies any of the conditions stated. So, without loss of generality, $\boldsymbol{A}\notin\mathbb C^d\boldsymbol{B}$. By \cite[Theorem 1.6, p. 201]{Singer}, $\boldsymbol{A}$ is orthogonal to $\C^d\boldsymbol{B}$ if and only if there exist $h$ functionals $f_1, f_2, \dots, f_h\in ((\C^n)^d)^*$ of unit norm with $1\leq h\leq 2d+1$, $h$ unit vectors $x_1, \dots, x_h\in {\C^n}$ and $t_1, \dots, t_h>0$ with $\sum\limits_{i=1}^ht_i =1$ such that \begin{equation}\label{1}f_i(\boldsymbol{A}x_i)=\|\boldsymbol{A}\| \text{ for all } 1\leq i\leq h\end{equation} and \begin{equation}\label{2}\sum\limits_{i=1}^nt_if_i(\boldsymbol{\lambda B}x_i) = 0 \text{ for all } \boldsymbol{\lambda}\in\C^d.\end{equation}
By the Riesz Representation Theorem, there exist unit vectors $y_1, \dots, y_h\in \Hil$ such that for $1\leq i\leq h$, $f_i(x) = \langle y_i| x\rangle$ for all $x\in\Hil$. So \eqref{1} is equivalent to the condition $\langle y_i|\boldsymbol{A}x_i\rangle =\|\boldsymbol{A}\|$. By the condition of equality in the Cauchy-Schwarz inequality, this is equivalent to $y_i = \dfrac{1}{\|\boldsymbol{A}\|}\boldsymbol{A}x_i$. So $\|\boldsymbol{A}x_i\|=\|\boldsymbol{A}\|$. Thus \eqref{2} is equivalent to $\sum\limits_{i=1}^ht_i \langle \boldsymbol{A}x_i| \boldsymbol{\lambda B}x_i\rangle =0$ for all $ \boldsymbol{\lambda}\in\C^d$, that is, for $1\leq k\leq d$, $\sum\limits_{i=1}^h t_i\langle A_kx_i|B_kx_i \rangle=0$.
\end{proof}
%Let $I$ be the identity operator on $\Hil$. A special case of Theorem \ref{newproof} for $\boldsymbol{B}=(I, \dots, I)$ was proved in \cite[Theorem 1]{2021} and it was used to show a distance formula for tuples of opeators was also mentioned. 

Let $\hc, \hk$ be Hilbert spaces. Let $\mathscr B(\hc, \hk)$ be the space of bounded operators from $\hc$ to $\hk$. The notation $\mathscr B(\hc)$ stands for $\mathscr B(\hc, \hc)$. In Theorem 2.8 of \cite{arxiv}, the following characterization is obtained. Let $A\in\mathscr B(\hc)$ be such that $\|A\|=1$, the set $\{x\in\hc:\|Ax\|=\|A\|\}$ is the unit ball of a finite dimensional subspace $\hc_1$ of $\hc$ and $\|A\|_{\hc_1^{\perp}}<\|A\|$. Then for any subspace $\mathcal W$ of $\mathscr B(\hc)$, $A$ is orthogonal to $\mathcal W$ if and only if there exist unit vectors $x_1, \dots, x_h\in\hc_1$ with $\|Ax_i\|=\|A\|$ for all $1\leq i\leq h$ and there exist $t_1, \dots, t_h>0$ with $\sum\limits_{i=1}^ht_i =1$ such that  $\sum\limits_{i=1}^h t_i\langle Ax_i | Bx_i\rangle =0$ for all $B\in\mathcal W$. Along the lines of the proof of Theorem \ref{newproof} above, we get the following generalization of this.
\begin{theorem}\label{bound} Let $A\in\mathscr B(\hc, \hk)$ be such that the set $\{x\in\hc:\|Ax\|=\|A\|\}$ is the unit ball of a finite dimensional subspace $\hc_1$ of $\hc$ and $\|A\|_{\hc_1^{\perp}}<\|A\|$. Then for any subspace $\mathcal W$ of $\mathscr B(\hc, \hk)$, $A$ is orthogonal to $\mathcal W$ if and only if there exist unit vectors $x_1, \dots, x_h\in\hc_1$ with $\|Ax_i\|=\|A\|$ for all $1\leq i\leq h$ and there exist $t_1, \dots, t_h>0$ with $\sum\limits_{i=1}^ht_i =1$ such that  $\sum\limits_{i=1}^h t_i\langle Ax_i | Bx_i\rangle =0$ for all $B\in\mathcal W$. Moreover, $1\leq h\leq 2\dim(\mathcal W)+1$.
\end{theorem}

Since the vectors $x_1, \dots, x_h$ can be chosen to be linearly independent, $\text{dim}(\hc)$ is also a bound on $h$.  Theorem 1 of \cite{2014} and Theorem 8.4 of \cite{Zietak2} are special cases of Theorem \ref{bound}. In both the papers, the bound on $h$ was shown to be $dim(\hc)$ and we have been able to find a better bound on $h$. A generalization of the above theorem without any condition on $A$ can be found in \cite[Theorem 1.3]{new}. When $\mathcal W$ is a one dimensional subspace, a characterization of orthogonality was first proved in \cite[Lemma 2.2]{magajna}. It was motivated by the proof of \cite[Lemma 9.14]{Davidson}.  An alternate proof of this can be found in \cite[Remark 3.1]{Bhatia}. For a detailed survey on orthogonality to subspaces and its applications, see \cite{2019, 2020} and the references therein.

\section{Remarks}\label{Remarks}

\textbf{Remark 1}. Let $X$ be a reflexive Banach space and $Y$ be a Banach space. Let $\mathscr K(X, Y)$ be the space of  compact operators from $X$ to $Y$ with the operator norm. For $x\in X$ and a subspace $W$ of $X$, let $\dist(x, W)=\inf\{\|x-w\|:w\in W\}$. Theorem \ref{newproof} also holds for $A\in\mathscr B(X, Y)$ such that $\dist(A, \mathscr K(X, Y))<\|A\|$. This can be seen from \cite[Lemma 3.1]{Wojcik} and the proof of Corollary \ref{101010}. An expression for the subdifferential set of the norm function in $\mathscr B(X, Y)$ for a reflexive Banach space $X$ was also obtained in \cite[Theorem 3.2]{Wojcik}.

%\textbf{Remark 2}. By the Haudorff-Toeplitz theorem and Theorem \ref{bound}, we get an alternative proof \cite[Theorem 1.1]{Bhatia} that gives a characterization of orthogonality of an element to one dimensional subspace in the space of matrices. A characterization of orthogonality of an element to one dimensional subspace in the space of operators on a Hilbert space was also obtained in \cite[Remark 3.1]{Bhatia} and \cite[Lemma2]{Radii}. We would like to mention that this result was first proved in \cite[Lemma 2.2]{magajna}, which seem to be missing from the literature on Birkhoff-James orthogonality. It was motivated by the proof of the \cite[Lemma 9.14]{Davidson}. And it is also interesting to note that \cite[Lemma 9.15]{Davidson} gives an alternate proof of  \cite[Theorem 4]{derivation}, and the idea of the proof of \cite[Theorem 1.2]{Bhatia} was along the same lines.

%\begin{remark} A natural thing to ask will be: Is it possible to find $h=1$ in the Theorem \ref{100000} or Theorem \ref{101010} or Theorem \ref{newproof}? And the answer is no, in general. A counterexample that $h$ may not be $1$ in Theorem \ref{newproof} was given in \cite{2021}. But in 
%\end{remark}

\textbf{Remark 2}. Birkhoff-James orthogonality is closely related to the notion of norm parallelism. In a normed space, an element $x$ is said to be \emph{norm parallel} to another element $y$ if there exists $\lambda\in\mathbb C$ such that $|\lambda|=1$ and $\|x+\lambda y\|=\|x\|+\|y\|$. Let $\boldsymbol{A}, \boldsymbol{B}\in\mathbb M_n(\C)^d$. Then by \cite[Theorem 2.4]{normparallel} and Theorem \ref{100000}, we get that $\boldsymbol{A}$ is norm parallel to $\boldsymbol{B}$ in the joint numerical radius if and only if there exists a unit vector $x\in\C^n$ such that $\big|\sum_{k=1}^d \overline{\bra x|B_kx\ket}\bra x|A_kx\ket\big|=\omega(\boldsymbol{A})\omega(\boldsymbol{B})$. The same characterization also holds for $\boldsymbol{A}, \boldsymbol{B}\in\mathscr B(\hc, \hc^d)$ for a Hilbert space $\hc$. The proof can be done along the lines of the proof of \cite[Theroem 2.2]{radiusparallel}.

\end{document}